# A Multi-Stage Hybrid CNN-Transformer Network for Automated Pediatric Lung Sound Classification

### v1

@@ Abstract
Automated analysis of lung sound auscultation is essential for monitoring respiratory health, especially in regions facing a shortage of skilled healthcare workers. While respiratory sound classification has been widely studied in adults, its application in pediatric populations, particularly in children aged $<$6 years, remains an underexplored area. The developmental changes in pediatric lungs considerably alter the acoustic properties of respiratory sounds, necessitating specialized classification approaches tailored to this age group. To address this, we propose a multistage hybrid CNN-Transformer framework that combines CNN-extracted features with an attention-based architecture to classify pediatric respiratory diseases using scalogram images from both full recordings and individual breath events. Our model achieved an overall score of 0.9039 in binary event classification and 0.8448 in multiclass event classification by employing class-wise focal loss to address data imbalance. At the recording level, the model attained scores of 0.720 for ternary and 0.571 for multiclass classification. These scores outperform the previous best models by 3.81% and 5.94%, respectively. This approach offers a promising solution for scalable pediatric respiratory disease diagnosis, especially in resource-limited settings.
Index Terms:
Lung auscultation sound, respiratory disease detection, transformer, continuous wavelet transform, scalogram, focal Loss.
@@ I Introduction
Shuvo et al. [10] (2020) Lightweight CNN Scalograms Pathological Condition ICBHI score 98.70% (6-class)
@@
Nguyen et al. [12] (2020) CNN Snapshot Ensemble Mel-spectrogram Breathing cycles Acc.: 78.4% (4-class), 83.7% (2-class)
Sudha et al. [13] (2021) Transfer learning Scalogram Breathing cycles Acc.: 92.50% (ICBHI) & 92.68% (RALE)
@@
Figure 1: A graphical overview of the proposed framework for lung sound analysis. The framework begins with the SPRsound dataset, followed by several pre-processing steps, including low-pass filtering, resampling, and normalization. The lung sound signals are then segmented based on labels and padded to a fixed length. These transformed signals are then fed into the proposed model, which performs classification tasks, including (i) normal vs. adventitious sounds (CAS, DAS) and (ii) categorization into specific respiratory sound events such as wheeze, crackle, and rhonchi.
@@ II-C Transformation into 2D images
[…]
Here, $a$ and $b$ represent the scale and translation parameters, respectively. The squared magnitude of the CWT coefficients $Z$ is referred to as the scalogram [40]. Both the 8.216-second (recorded) and 3-second (segmented) lung sound samples are analyzed in the wavelet domain using the Morse analytic wavelet as the mother wavelet, and the scalogram images of $\times 224$ 224 resolution are produced. The time-bandwidth product and symmetry parameter for the wavelet are set to 60 and 3, respectively. The range of energy for the wavelet in both time and frequency determines the minimum and maximum scales, which are automatically chosen using 10 voices per octave [43].
@@ III-A Feature Extraction Block
The feature extraction block utilizes MobileNetV2 [44], a lightweight CNN optimized for efficient feature extraction. MobileNetV2 employs depthwise separable convolutions, significantly reducing computational complexity while retaining rich feature representations. The input to the model consists of 2D scalogram representations of lung sound signals with dimensions 224×224×3224$\times$ 224$\times$ 3224 × 224 × 3. The first convolutional layer applies a $\times 3$ 3 kernel with 32 filters, followed by a series of inverted residual blocks with linear bottlenecks to enhance feature expressiveness. The depthwise separable convolution is mathematically formulated as follows:
@@
This operation significantly reduces the number of parameters compared to standard convolutions. The final output from MobileNetV2 is a feature map of size 7×7×12807$\times$ 7$\times$ 12807 × 7 × 1280, which is further refined in the next stage.
@@ IV-B Evaluation Criteria
The training audio files (both for recording level and event level classification) are split into two sets with an approximate proportion of 90:1090:1090 : 10 split for creating training and validation sets. The provided testing set defined by the challenge was used to evaluate the model.
@@ IV-C Hyperparameters
The proposed model is constructed using TensorFlow backend and trained using NVidia K80 GPUs. In this study, a batch size of 128128128 has been selected for all the Tasks. Optimization is performed using the Adam optimizer with a learning rate of 0.001.
@@ VII Conclusion
This study addressed the precise classification of pediatric respiratory diseases at both the event and recording levels. For that, we proposed a multi-stage hybrid CNN-Transformer model for classifying lung sounds, which effectively combines MobileNet-V2 for efficient feature extraction with transformer-based self-attention for enhanced feature refinement. The model demonstrated superior performance across event-level and recording-level tasks and has achieved a significantly high overall score of 0.90390.90390.9039 and 0.84480.84480.8448 for Task 1-1 and Task 1-2, respectively, in breath event classification. Similarly, the model outperformed all the SOTA systems at Task 2-2 with a score of 0.57100.57100.5710. However, the model has limitations, including reliance on a relatively small dataset and high computational demands.

### v2

@@ Abstract
Background: Automated analysis of lung sound auscultation is essential for monitoring respiratory health, particularly in regions with a shortage of skilled healthcare workers. Although respiratory sound classification has been widely studied in adults, its application in pediatric populations, especially in children under six years of age remains underexplored. Developmental changes in pediatric lungs substantially modify the acoustic properties of respiratory sounds, requiring classification approaches tailored specifically to this age group.
Methods: To address this challenge, we propose a multistage hybrid CNN–Transformer framework that integrates CNN-extracted features with an attention-based architecture for pediatric respiratory disease classification. Scalogram images were generated from both full recordings and individual breath events to capture multi-resolution representations of respiratory sounds. To mitigate class imbalance, class-wise focal loss was applied during model training.
Results: The proposed model achieved an overall score of 0.9039 in binary event classification and 0.8448 in multiclass event classification. At the recording level, the model obtained scores of 0.720 for ternary classification and 0.571 for multiclass classification. These results outperform the previous best-performing models by 3.81% and 5.94%, respectively.
Conclusion: Our findings demonstrate that the proposed hybrid CNN–Transformer framework effectively captures the unique acoustic features of pediatric lung sounds. This approach offers a scalable and robust solution for automated pediatric respiratory disease diagnosis, with strong potential for deployment in resource-limited healthcare settings.
@@ I Introduction
Shuvo et al. [10] (2020) Lightweight CNN Scalograms Pathological Condition ICBHI score 98.70%
@@
Nguyen et al. [12] (2020) CNN Snapshot Ensemble Mel-spectrogram Breathing cycles Acc.: 78.4% (4-class) 83.7% (2-class)
Sudha et al. [13] (2021) Transfer learning Scalogram Breathing cycles Acc.: 92.50% (ICBHI) 92.68% (RALE)
@@
Figure 1: A graphical overview of the proposed framework for lung sound analysis. The framework begins with the SPRsound dataset, followed by several pre-processing steps, including low-pass filtering, resampling, and normalization. The lung sound signals are then segmented based on labels and padded to a fixed length. These transformed signals are then fed into the proposed model, which performs classification tasks
@@ II-C Transformation into 2D images
$Z(a,b)=\frac{1}{\sqrt{b}}\int x(t)g\left(\frac{t-a}{b}\right)dt$ (1)
Here, $a$ and $b$ represent the scale and translation parameters, respectively. The squared magnitude of the CWT coefficients $Z$ is referred to as the scalogram [40]. Both the 8.216-second (recorded) and 3-second (segmented) lung sound samples are analyzed in the wavelet domain using the Morse analytic wavelet as the mother wavelet, and the scalogram images of $224\times 224$ resolution are produced. The time-bandwidth product and symmetry parameter for the wavelet are set to 60 and 3, respectively. The range of energy for the wavelet in both time and frequency determines the minimum and maximum scales, which are automatically chosen using 10 voices per octave [43].
@@ III-A Feature Extraction Block
The feature extraction block utilizes MobileNetV2 [44], a lightweight CNN optimized for efficient feature extraction. MobileNetV2 employs depthwise separable convolutions, significantly reducing computational complexity while retaining rich feature representations. The input to the model consists of 2D scalogram representations of lung sound signals with dimensions $224\times 224\times 3$. The first convolutional layer applies a $3\times 3$ kernel with 32 filters, followed by a series of inverted residual blocks with linear bottlenecks to enhance feature expressiveness. The depthwise separable convolution is mathematically formulated as follows:
@@
This operation significantly reduces the number of parameters compared to standard convolutions. The final output from MobileNetV2 is a feature map of size $7\times 7\times 1280$, which is further refined in the next stage.
@@ IV-B Evaluation Criteria
The training audio files (both for recording level and event level classification) are split into two sets with an approximate proportion of $90:10$ split for creating training and validation sets. The provided testing set defined by the challenge was used to evaluate the model.
@@ IV-C Hyperparameters
The proposed model is constructed using TensorFlow backend and trained using NVidia K80 GPUs. In this study, a batch size of $128$ has been selected for all the Tasks. Optimization is performed using the Adam optimizer with a learning rate of 0.001.
@@ VII Conclusion
This study addressed the precise classification of pediatric respiratory diseases at both the event and recording levels. For that, we proposed a multi-stage hybrid CNN-Transformer model for classifying lung sounds, which effectively combines MobileNet-V2 for efficient feature extraction with transformer-based self-attention for enhanced feature refinement. The model demonstrated superior performance across event-level and recording-level tasks and has achieved a significantly high overall score of $0.9039$ and $0.8448$ for Task 1-1 and Task 1-2, respectively, in breath event classification. Similarly, the model outperformed all the SOTA systems at Task 2-2 with a score of $0.5710$. However, the model has limitations, including reliance on a relatively small dataset and high computational demands.
CRediT authorship contribution statement
Samiul Based Shuvo: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Resources, Methodology, Formal analysis, Data curation, Conceptualization. Taufiq Hasan: Writing – review & editing, Supervision, Project administration, Investigation.